What is the value of the discriminant?
X(squared) - 3x - 12 = 0

The teacher collected the textbooks today, the final is tomorrow.

Answer by fractalier(6550)

For x^2 - 3x - 12 = 0
we examine the discriminant as b^2 - 4ac...here that is
(-3)^2 - 4(1)(-12) =
9 + 48 =
57